# Tight PAC-Bayesian Risk Certificates for Contrastive Learning

### v2

@@ 1 Introduction
A key driving force behind the rapid advances in foundation models is the availability and exploitation of massive amounts of unlabeled data. Broadly, one learns meaningful representations from unlabeled data, reducing the demand for labeled samples when training (downstream) predictive models. In recent years, there has been a strong focus on self-supervised approaches to representation learning, which learn neural network-based embedding maps from carefully constructed augmentations of unlabeled data, such as image cropping, rotations, color distortion, Gaussian blur, etc. [3, 7, 9, 15]. Contrastive representation learning is a popular form of self-supervised learning where one aims to learn a mapping of the data to a Euclidean space such that semantically similar data, obtained via augmentations, are embedded closer than independent samples [45, 19, 24]. This technique gained widespread attention with the introduction of SimCLR, an abbreviation for simple framework for contrastive learning of representations [7].
@@
The SimCLR framework employs a carefully designed contrastive loss to maximize the similarity between the representations of augmented views of the same sample while minimizing the similarity between the representations from different samples [39, 7]. Although SimCLR remains one of the most practically used contrastive models, theoretical analysis of SimCLR’s performance and generalization abilities is still limited [4, 32]. The study of generalization error in self-supervised models is mostly based on two distinct frameworks [2, 17], both introduced in the context of contrastive learning. The contrastive unsupervised representation learning (CURL) framework, introduced by Arora et al. [2], assumes access to tuples $z_{1},\ldots,z_{n}$, where each $z=(x,x^{+},x^{-}_{1},\ldots,x^{-}_{k})$. The underlying statistical model considers data from a mixture of $k$ (class) distributions, and within each $z$, it is assumed that $x$ and $x^{+}$ are independent samples from the same class, while $x^{-}_{1},\ldots,x^{-}_{k}$ are $k$ i.i.d. samples from the mixture model, independent of $x,x^{+}$. Arora et al. [2] derive Rademacher complexity-based generalization error (risk) bounds for representation learning using bound contrastive losses. Furthermore, the above statistical model for augmented data $z$ allows one to extend the bounds to derive a risk certificate for downstream classification using the mean classifier [2]. Subsequently, improved risk certificates for both contrastive learning and downstream classification were derived in [26] using empirical covering numbers, and also in [32] using a PAC-Bayesian analysis or in [4] — the latter bounds being specifically obtained for the N-pair (or InfoNCE) loss used in SimCLR [39, 7]. The CURL framework has been used to derive risk bounds for non-contrastive models [5] and adversarial contrastive models [48], although both rely on Rademacher bounds. PAC-Bayes bounds have been instead stated in the context of meta-learning [12].
@@
The independence of samples assumed in the CURL framework is quite impractical. For instance, in the SimCLR framework, the positive samples $x,x^{+}$ are augmented views of the same data instances, for example, random rotation or cropping of the same image. In addition, practical implementations of SimCLR do not generate independent negative samples. The loss is computed over batches of positive pairs $(x_{i},x_{i}^{+})_{i=1,\ldots,m}$ and for each $x_{i}$ in a batch, all $x_{j},x_{j}^{+},j\neq i$ are used as negative samples, thereby inducing strong dependence across tuples $z_{1},\ldots,z_{n}$ and making classical PAC-Bayes analysis inapplicable. Nozawa et al. [32] derive $f$-divergence-based PAC-Bayes bounds to account for potential dependence of the positive/negative samples within each tuple $z$, but fail to account for dependence across tuples $z_{1},\ldots,z_{n}$.
@@
Downstream Risk Many --ve samples Data augmentation Temperature scaling
@@
The paper is organized as follows. We provide background on the SimCLR framework and PAC-Bayes theory in section 2. We then present our main contribution in section 3. In section 4, we evaluate our results by conducting experiments on MNIST and CIFAR-10, and a discussion follows in section 5.
@@ 2.1 SimCLR framework
[…]
${\widehat{L}}_{S}(f)=\frac{1}{n}\sum_{i=1}^{n}\frac{\ell_{\operatorname{cont}}(x_{i},x_{i}^{+},\bigcup_{j\neq i}\{x_{j}\})+\ell_{\operatorname{cont}}(x_{i}^{+},x_{i},\bigcup_{j\neq i}\{x_{j}^{+}\})}{2}=\frac{1}{n}\sum_{i=1}^{n}\ell(x_{i},x_{i},X_{i}^{-}),$ (1)
where $\ell(x_{i},x_{i},X_{i}^{-})$ denotes the above symmetrization of $\ell_{\operatorname{cont}}$.
@@
The SimCLR loss presented in eq. 1 differs slightly from the original loss, as it considers a subset of negative samples $X_{i}^{-}$ instead of the full set $X^{-}$ in each $\ell_{\operatorname{cont}}$. This approach still uses all available negative samples while simplifying the theoretical derivations. Importantly, any results derived for this simplified SimCLR loss can be straightforwardly extended to the original SimCLR loss [7].
@@
[…]
@@
$R(f)$ evaluates the representation $f$ in terms of how often embeds positive pairs further than negative samples. Since the SimCLR loss can be seen as a surrogate loss for the contrastive zero-one risk, risk certificates for the contrastive zero-one risk prove to be valuable [32].
@@ 2.2 PAC-Bayes Theory
[…]
@@ 3.1 Risk Certificates for Contrastive Loss
Since the SimCLR loss is computed over a batch, it induces dependence across the summand $\ell$ in (1) through $X^{-}$, violating the assumption required in the PAC-Bayes-kl bound (theorem 1). Since the dataset $S$ is partitioned into $p=\frac{n}{m}$ i.i.d. batches, one simple workaround is to compute the PAC-Bayes bound over these i.i.d. batches instead of the usual $n$ i.i.d. samples. For a detailed overview of the bounds, please refer to the supplementary materials. However, this weakens the risk certificate when there are few large batches. Alternatively, Nozawa et al. [32] suggest using $f$-divergence to address the non-i.i.d. characteristics, but their proposed bounds are vacuous. In this section, we present our first contribution: two PAC-Bayesian risk certificates specifically tailored to the SimCLR framework.
@@ 3.1.1 First PAC-Bayes certificate
We can indeed prove that the SimCLR loss (1) satisfies the bounded difference assumption when one views the loss as a function of $\bar{x}_{1},\ldots,\bar{x}_{n}\sim_{iid}\mathcal{D}_{\mathcal{X}}$ that generates the positive pairs (see lemma 1). This assumption allows us to use McDiarmid’s inequality [30]. Although one cannot directly incorporate this assumption into the proof of the PAC-Bayes-kl bound, it is possible to include McDiarmid’s inequality in the proof of the PAC-Bayes bound presented in McAllester’s work [28]. We present our result in the following theorem.
@@
Let $\delta\in(0,1)$ be a confidence parameter and $C=4+(m-1)\log\frac{(m-1)+e^{2}}{m}$. With probability at least $1-\delta$ over dataset $S\sim\mathcal{S}^{n}$, for all $Q$:
@@
One can observe that theorem 3 holds for the original SimCLR loss as it satisfies the bounded difference assumption for $C=\frac{4}{\tau}+2(m-1)\log\frac{2(m-1)+e^{\frac{2}{\tau}}}{2m-1}$. To prove this, every occurrence of $m-1$ can be replaced by $2(m-1)$ in the proof presented below.
The rest of the subsection proves theorem 3. We first show that the SimCLR loss satisfies the bounded difference assumption.
@@
Proof of lemma 1.
Let $i\in[1,n]$. We aim to bound the following quantity:
$\Delta\phi(\bar{x}_{i})=\phi\left(\bar{x}_{1},\ldots,\bar{x}_{i},\ldots,\bar{x}_{n}\right)-\phi\left(\bar{x}_{1},\ldots,\bar{x}_{i}^{\prime},\ldots,\bar{x}_{n}\right).$
Without loss of generality, we assume that $i\in\{1,\dots,m\}$, meaning $\bar{x}_{i}$ belongs to the first batch. The quantity $\Delta\phi(\bar{x}_{i})$ can be split into two non-null terms:
$\Delta\phi(\bar{x}_{i})=\frac{1}{n}\left[\delta_{i}(\bar{x}_{i})+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{m}\delta_{j}(\bar{x}_{i})\right],$
@@
$\delta_{i}(\bar{x}_{i})$ $:=\ell(x_{i},x_{i}^{+},X_{i}^{-})-\ell(x_{i}^{\prime},x_{i}^{\prime+},X_{i}^{-}),$
$\delta_{j}(\bar{x}_{i})$ $:=\ell((x_{j},x_{j}^{+}),X_{j}^{-})-\ell((x_{j},x_{j}^{+}),\widetilde{X}_{j}^{-}),$
and $\widetilde{X}_{j}^{-}$ is the set of negative samples perturbed with $x_{i}^{\prime}$. Since $\delta_{i}(\bar{x}_{i})$ can be written as an average of two terms that play a symmetric role, upper-bounding $\delta_{i}(\bar{x}_{i})$ reduces to upper-bounding $\delta_{i}(x_{i})=\ell_{cont}((x_{i},x_{i}^{+}),X_{i})-\ell_{cont}((x_{i}^{\prime},x_{i}^{\prime+}),X_{i}),$ where, for convenience, we denote $X_{i}=\bigcup_{j\neq i}\{x_{j}\}$. Using the notation $S(x,X):=\sum_{x^{\prime}\in X}\operatorname{sim}(x,x^{\prime})$, we have:
$\delta_{i}(x_{i})=\log\frac{\operatorname{sim}(x_{i}^{\prime},x_{i}^{\prime+})}{\operatorname{sim}(x_{i},x_{i}^{+})}+\log\frac{\operatorname{sim}(x_{i},x_{i}^{+})+S(x_{i},X_{i})}{\operatorname{sim}(x_{i}^{\prime},x_{i}^{\prime+})+S(x_{i}^{\prime},X_{i})}\leq\frac{4}{\tau},$
as $\forall x,y,\ e^{-\frac{1}{\tau}}\leq\operatorname{sim}(x,y)\leq e^{\frac{1}{\tau}}$. Similarly, the term $\delta_{j}(\bar{x}_{i})$ can also be written as an average of two terms. It suffices to upper-bound $\delta_{j}(x_{i}):=\ell_{cont}((x_{j},x_{j}^{+}),X_{j})-\ell_{cont}((x_{j},x_{j}^{+}),\widetilde{X}_{j}),$ where $\widetilde{X}_{j}^{-}$ is the set of $X_{j}$ perturbed with $x_{i}^{\prime}$.. This can be rewritten as:
$\delta_{j}(x_{i})=\log\frac{\kappa+\operatorname{sim}(x_{j},x_{i})}{\kappa+\operatorname{sim}(x_{j},x_{i}^{\prime})}\leq\frac{(m-1)+e^{\frac{2}{\tau}}}{m},$
using $(m-1)e^{-\frac{1}{\tau}}\leq\kappa\leq(m-1)e^{\frac{1}{\tau}}$. Combining these results, we obtain $|\delta_{i}(\bar{x}_{i})|\leq\frac{4}{\tau}$ and $|\delta_{j}(\bar{x}_{i})|\leq\log\frac{(m-1)+e^{\frac{2}{\tau}}}{m}$. This gives:
$\Delta\phi(\bar{x}_{i})\leq\frac{1}{n}\left(\frac{4}{\tau}+(m-1)\log\frac{(m-1)+e^{\frac{2}{\tau}}}{m}\right)=\frac{C}{n}.$
@@
The rest of the proof of theorem 3 follows the different steps of the proof presented in McAllester [28], while incorporating the bounded difference assumption. The first step involves showing that we can bound the quantity $\underset{S\sim\mathcal{S}^{n}}{\mathbb{E}}\left[e^{(n-1)h\left(L_{S}(f)-L(f)\right)}\right]$ for a certain non-negative and convex function $h$ (this corresponds to lemmas 5 and 6 in [28]). lemma 3 below modifies lemma 6 in [28] since we cannot use Hoeffding’s inequality, whereas lemma 2 is a minor modification of lemma 5 in [28] that allows us to directly apply McDiarmid’s inequality to prove lemma 3.
@@
The bounded difference assumption allows us to derive the following McDiarmid’s inequality: for $\varepsilon>0$, $\mathbb{P}_{S}\left(|L(f)-L_{S}(f)|\geq\varepsilon\right)\leq 2\exp\left(-\frac{2\varepsilon^{2}}{\sum_{i=1}^{n}c_{i}^{2}}\right)$. Using the change of variable $X=\frac{\sqrt{2}}{C}(L_{S}(f)-L(f))$, we observe that the McDiarmid’s inequality gives: for $x>0$, $\mathbb{P}\left(\frac{\sqrt{2}}{C}|X|\geq x\right)\leq 2\exp\left(-nx^{2}\right)$. Invoking lemma 2 , we obtain $\mathbb{E}_{S}\left[e^{(n-1)h(X)}\right]\leq 2n$ where $h:x\mapsto\frac{2x^{2}}{C^{2}}$. ∎
Given the bound in lemma 3, one can derive the PAC-Bayes bound as shown by McAllester [28] (lemma 7–8). We adapt those results in the following statement.
@@
Assume for a fixed $f\sim P$, $\underset{S\sim\mathcal{S}^{n}}{\mathbb{E}}\left[e^{(n-1)h\left(L_{S}(f)-L(f)\right)}\right]\leq 2n$. This implies the bound
$\mathbb{E}_{S}\left[\mathbb{E}_{f\sim P}\left[e^{(n-1)h\left(L_{S}(f)-L(f)\right)}\right]\right]\leq 2n.$
Applying Markov’s inequality, we obtain that with probability at least $1-\delta$ over $S$,
$\mathbb{E}_{f\sim P}\left[e^{(n-1)h\left(L_{S}(f)-L(f)\right)}\right]\leq\frac{2n}{\delta}$
Next, we use a shift of measure: $\mathbb{E}_{f\sim Q}\left[(n-1)h\left(L_{S}(f)-L(f)\right)\right]\leq\mathrm{KL}(Q\|P)+\log\mathbb{E}_{f\sim P}\left[e^{(n-1)h\left(L_{S}(f)-L(f)\right)}\right]$. Combining the previous results, with probability at least $1-\delta$ over dataset $S$:
@@
Rewriting the statement of lemma 4 shows that, with probability at least $1-\delta$ over i.i.d. dataset $S$, $L(Q)\leq L_{S}(Q)+C\sqrt{\frac{\mathrm{KL}(Q\|P)+\log\frac{2n}{\delta}}{2(n-1)}}$ for all $Q$, completing the proof of theorem 3.

### v4

@@ 1 Introduction
A key driving force behind the rapid advances in foundation models is the availability and exploitation of massive amounts of unlabeled data. Broadly, one learns meaningful representations from unlabeled data, reducing the demand for labeled samples when training (downstream) predictive models. In recent years, there has been a strong focus on self-supervised approaches to representation learning, which learn neural network-based embedding maps from carefully constructed augmentations of unlabeled data, such as image cropping, rotations, color distortion, Gaussian blur, etc. [2, 7, 9, 17].
Contrastive representation learning is a popular form of self-supervised learning where one aims to extract meaningful features from unlabeled data by distinguishing between similar samples, obtained via augmentations, and dissimilar samples [54, 22, 28]. It has shown strong empirical performance across a wide range of of applications [24], such as image recognition [7, 57], natural language processing [14], audio processing [12], and graph processing [55, 42]. The core idea is to learn an embedding space where similar (positive) samples are pulled closer together, while dissimilar (negative) samples are pushed apart. This is typically accomplished using a contrastive loss function applied to features extracted by an encoder. A standard contrastive learning pipeline begins with data augmentation, where two augmented views of the same sample are generated to form a positive pair, while other samples in the batch serve as negatives. These augmented views are passed through an encoder network, optionally followed by a projection head. The network is then trained using a contrastive loss computed on these feature representations. Central to many contrastive frameworks is the InfoNCE loss [48, 18, 37], which uses a softmax function over pairwise feature similarities, often incorporating feature normalization and temperature scaling. This loss is employed in popular frameworks such as SimCLR [7], MoCo [22], and CLIP [43], where negative samples play a critical role. In contrast, other methods like Barlow Twins [56] and VICReg [4] eliminate the need for negative sampling by focusing on constraining the cross-correlation or cross-covariance matrix, while still ensuring alignment between positive pairs.
@@
In this work, we focus on the SimCLR framework (i.e., simple framework for contrastive learning of representations), which has gained widespread attention and demonstrated strong performance in a variety of downstream tasks[7]. Although SimCLR remains one of the most practically used contrastive models, theoretical analysis of SimCLR’s performance and generalization abilities is still limited [3, 36]. The study of generalization error in self-supervised models is mostly based on two distinct frameworks [44, 20], both introduced in the context of contrastive learning. The contrastive unsupervised representation learning (CURL) framework, introduced by Arora et al. [44], assumes access to tuples $z_{1},\ldots,z_{n}$, where each $z=(x,x^{+},x^{-}_{1},\ldots,x^{-}_{k})$. The underlying statistical model considers data from a mixture of $k$ (class) distributions, and within each $z$, it is assumed that $x$ and $x^{+}$ are independent samples from the same class, while $x^{-}_{1},\ldots,x^{-}_{k}$ are $k$ i.i.d. samples from the mixture model, independent of $x,x^{+}$. Arora et al. [44] derive Rademacher complexity-based generalization error (risk) bounds for representation learning using bound contrastive losses. Furthermore, the above statistical model for augmented data $z$ allows one to extend the bounds to derive a risk certificate for downstream classification using the mean classifier [44]. Subsequently, improved risk certificates for both contrastive learning and downstream classification were derived in [30] using empirical covering numbers, and also in [36] using a PAC-Bayesian analysis or in [3] — the latter bounds being specifically obtained for the N-pair (or InfoNCE) loss used in SimCLR [48, 7]. The CURL framework has been used to derive risk bounds for non-contrastive models [5] and adversarial contrastive models [59], although both rely on Rademacher bounds. PAC-Bayes bounds have been instead stated in the context of meta-learning [13].
@@
The independence of samples assumed in the CURL framework is quite impractical. For instance, in the SimCLR framework, the positive samples $x,x^{+}$ are augmented views of the same data instances, for example, random rotation or cropping of the same image. In addition, practical implementations of SimCLR do not generate independent negative samples. The loss is computed over batches of positive pairs $(x_{i},x_{i}^{+})_{i=1,\ldots,m}$ and for each $x_{i}$ in a batch, all $x_{j},x_{j}^{+},j\neq i$ are used as negative samples, thereby inducing strong dependence across tuples $z_{1},\ldots,z_{n}$ and making classical PAC-Bayes analysis inapplicable. Nozawa et al. [36] derive $f$-divergence-based PAC-Bayes bounds to account for potential dependence across tuples $z_{1},\ldots,z_{n}$; however, their resulting bounds are vacuous in practice.
@@
Downstream Risk Many -ve samples Data augmentation Temperature scaling
@@
The paper is organized as follows. We provide background on the SimCLR framework and PAC-Bayes theory in section˜2. We then present our main contribution in section˜3. In section˜4, we evaluate our results by conducting experiments on MNIST and CIFAR-10, and a discussion follows in section˜5.
@@ 2.1 SimCLR framework
In SimCLR, the similarity between the learned representations of two instances $f(x)$ and $f(x^{\prime})$ is typically defined in terms of their cosine similarity, $\frac{f(x)^{\top}f(x^{\prime})}{\|f(x)\|\;\|f(x^{\prime})\|}$, where $\|\cdot\|$ is the Euclidean norm. Since the feature representations are normalized, we express the similarity between two representations as $f(x)^{\top}f(x^{\prime})$. Let $\tau\in(0,\infty)$ be a temperature parameter. To improve readability, we use the notation $\operatorname{sim}(x,x^{\prime})=\exp\left(\frac{f(x)^{\top}f(x^{\prime})}{\tau}\right)$ and define $\ell_{\operatorname{cont}}(x,x^{+},X)=-\log\frac{\operatorname{sim}(x,x^{+})}{\operatorname{sim}(x,x^{+})+\sum_{x^{\prime}\in X}\operatorname{sim}(x,x^{\prime})}$ . The empirical SimCLR loss over the dataset $S\sim\mathcal{S}^{n}$, denoted ${\widehat{L}}_{S}(f)$, is given by [7]:
$\frac{1}{n}\sum_{i=1}^{n}\frac{\ell_{\operatorname{cont}}(x_{i},x_{i}^{+},\bigcup_{j\neq i}\{x_{j}\})+\ell_{\operatorname{cont}}(x_{i}^{+},x_{i},\bigcup_{j\neq i}\{x_{j}^{+}\})}{2}=\frac{1}{n}\sum_{i=1}^{n}\ell(x_{i},x_{i}^{+},X_{i}^{-}),$ (1)
where $\ell(x_{i},x_{i}^{+},X_{i}^{-})$ denotes the above symmetrization of $\ell_{\operatorname{cont}}$.
@@
The SimCLR loss presented in eq.˜1 differs slightly from the original loss from [7], as it considers a subset of negative samples $X_{i}^{-}$ instead of the full set $X^{-}$ in each $\ell_{\operatorname{cont}}$. This approach still uses all available negative samples while simplifying the theoretical derivations. Importantly, any results derived for this simplified SimCLR loss can be straightforwardly extended to the original SimCLR loss [7]. A detailed expression of the original loss is provided in the supplementary materials.
@@
$\texttt{top1}=1-R_{\text{top1}}(f,W),\quad\text{where }R_{\text{top1}}(f,W)=\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\mathbb{I}_{\{f(x)^{\top}w_{y}<\max\limits_{c\neq y}f(x)^{\top}w_{c}\}}\right].$ (4)
@@
$R(f)$ evaluates the representation $f$ in terms of how often it embeds positive pairs further than negative samples. Since the SimCLR loss can be seen as a surrogate loss for the contrastive zero-one risk, risk certificates for the contrastive zero-one risk prove to be valuable [36].
@@ 2.2 PAC-Bayes Theory
Let $P$ denote a prior distribution and $Q$ a posterior distribution over the parameter space $\mathcal{W}$. In PAC-Bayes theory, the distance between the prior and posterior distributions is often quantified using the Kullback-Leibler (KL) divergence, defined as $\mathrm{KL}\left(Q\|Q^{\prime}\right)=\int\limits_{\mathcal{W}}\log\left(\frac{dQ}{dQ^{\prime}}\right)dQ.$ The binary KL divergence is $\mathrm{kl}\left(q\|q^{\prime}\right)=q\log\left(\frac{q}{q^{\prime}}\right)+(1-q)\log\left(\frac{1-q}{1-q^{\prime}}\right)$, where $q,q^{\prime}\in[0,1]$—this measure quantifies the divergence between two Bernoulli distributions with parameters $q,q^{\prime}$. Let $\mathcal{Z}$ denote an example space, $\mathcal{D}_{\mathcal{Z}}$ the distribution over $\mathcal{Z}$, and $\ell_{w}:\mathcal{Z}\rightarrow[0,1]$ a loss function parameterized by $w\in\mathcal{W}$. The risk $L:\mathcal{W}\rightarrow[0,1]$ is defined as $L(w)=\underset{z\sim\mathcal{D}_{\mathcal{Z}}}{\mathbb{E}}[\ell_{w}(z)].$ Here, $L(w)$ represents the expected value of $\ell_{w}(z)$ under the distribution $\mathcal{D}_{\mathcal{Z}}$. Let $n$ be an integer, and the empirical risk for a dataset $S=(z_{1},\ldots,z_{n})\in\mathcal{Z}^{n}$ is defined as $\widehat{L}_{S}(w)=\frac{1}{n}\sum_{i=1}^{n}\ell_{w}(z_{i}).$ Here, $\widehat{L}_{S}(w)$ computes the average loss $\ell_{w}(z_{i})$ over the dataset $S$.
@@ 3.1 Risk Certificates for Contrastive Loss
Since the SimCLR loss is computed over a batch, it induces dependence across the summand $\ell$ in (1) through $X^{-}$, violating the assumption required in the PAC-Bayes-kl bound (theorem˜1). Since the dataset $S$ is partitioned into $p=\frac{n}{m}$ i.i.d. batches, one simple workaround is to compute the PAC-Bayes bound over these i.i.d. batches instead of the usual $n$ i.i.d. samples. For a detailed overview of the bounds, please refer to the supplementary materials. However, this weakens the risk certificate when there are few large batches. Alternatively, Nozawa et al. [36] suggest using $f$-divergence to address the non-i.i.d. characteristics, but their proposed bounds are vacuous. In this section, we present our first contribution: two PAC-Bayesian risk certificates specifically tailored to the SimCLR framework.
@@ 3.1.1 First PAC-Bayes certificate
We can indeed prove that the SimCLR loss (1) satisfies the bounded difference assumption when one views the loss as a function of $\bar{x}_{1},\ldots,\bar{x}_{n}\sim_{iid}\mathcal{D}_{\mathcal{X}}$ that generates the positive pairs (see lemma˜1). This assumption allows us to use McDiarmid’s inequality [34]. Although one cannot directly incorporate this assumption into the proof of the PAC-Bayes-kl bound, it is possible to include McDiarmid’s inequality in the proof of the PAC-Bayes bound presented in McAllester’s work [32]. We present our result in the following theorem.
@@
Let $\delta\in(0,1)$ be a confidence parameter and $C=4+(m-1)\log\frac{(m-1)+e^{2/\tau}}{m}$. With probability at least $1-\delta$ over dataset $S\sim\mathcal{S}^{n}$, for all $Q$:
@@
One can observe that theorem˜3 holds for the original SimCLR loss as it satisfies the bounded difference assumption for $C=\frac{4}{\tau}+(m-1)\log\frac{2m-3+2e^{2/\tau}}{2m-1}$. The detailed proof is provided in the supplementary materials.
The rest of the subsection proves theorem˜3. We first show that the SimCLR loss satisfies the bounded difference assumption.
@@
Proof of lemma˜1.
Let $i\in[1,n]$ and $\bar{S}=\{\bar{x}_{i}\}_{i=1}^{n}$. We aim to bound the following quantity:
$\Delta\phi_{\bar{S}}(\bar{x}_{i})=\phi\left(\bar{x}_{1},\ldots,\bar{x}_{i},\ldots,\bar{x}_{n}\right)-\phi\left(\bar{x}_{1},\ldots,\bar{x}_{i}^{\prime},\ldots,\bar{x}_{n}\right),$
where the notation emphasizes that the difference $\Delta\phi_{\bar{S}}(\bar{x}_{i})$ arises from perturbing the $i$-th argument of the function $\phi$, while all other arguments $\bar{x}_{j}$ for $j\neq i$ are held fixed.
Without loss of generality, we assume that $i\in\{1,\dots,m\}$, meaning $\bar{x}_{i}$ belongs to the first batch $\bar{S}_{m}=\{\bar{x}_{i}\}_{i=1}^{m}$. The quantity $\Delta\phi_{\bar{S}}(\bar{x}_{i})$ can be split into two non-null terms:
$\Delta\phi_{\bar{S}}(\bar{x}_{i})=\frac{1}{n}\left[\delta^{i}_{\bar{S}_{m}}(\bar{x}_{i})+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{m}\delta^{j}_{\bar{S}_{m}}(\bar{x}_{i})\right],$
@@
$\delta^{i}_{\bar{S}_{m}}(\bar{x}_{i})$ $=\ell(x_{i},x_{i}^{+},X_{i}^{-})-\ell(x_{i}^{\prime},x_{i}^{\prime+},X_{i}^{-}),$
$\delta^{j}_{\bar{S}_{m}}(\bar{x}_{i})$ $=\ell(x_{j},x_{j}^{+},X_{j}^{-})-\ell(x_{j},x_{j}^{+},\widetilde{X}_{j}^{-}),$
and $\widetilde{X}_{j}^{-}$ is the set of negative samples perturbed with $x_{i}^{\prime}$. Since the loss $\ell$ can be written as an average of two terms that play a symmetric role, upper-bounding $\delta^{i}_{\bar{S}_{m}}(\bar{x}_{i})$ reduces to upper-bounding $\ell_{cont}(x_{i},x_{i}^{+},X_{i})-\ell_{cont}(x_{i}^{\prime},x_{i}^{\prime+},X_{i}),$ where, for convenience, we denote $X_{i}=\bigcup_{j\neq i}\{x_{j}\}$. Using the notation $S(x,X):=\sum_{x^{\prime}\in X}\operatorname{sim}(x,x^{\prime})$, we have:
$|\ell_{cont}(x_{i},x_{i}^{+},X_{i})-\ell_{cont}(x_{i}^{\prime},x_{i}^{\prime+},X_{i})|=\left|\log\frac{\operatorname{sim}(x_{i}^{\prime},x_{i}^{\prime+})}{\operatorname{sim}(x_{i},x_{i}^{+})}+\log\frac{\operatorname{sim}(x_{i},x_{i}^{+})+S(x_{i},X_{i})}{\operatorname{sim}(x_{i}^{\prime},x_{i}^{\prime+})+S(x_{i}^{\prime},X_{i})}\right|\leq\frac{4}{\tau},$
as $\forall x,y,\ e^{-\frac{1}{\tau}}\leq\operatorname{sim}(x,y)\leq e^{\frac{1}{\tau}}$. Similarly, the term $\delta_{j}(\bar{x}_{i})$ can also be written as an average of two terms. It suffices to upper-bound $\ell_{cont}(x_{j},x_{j}^{+},X_{j})-\ell_{cont}(x_{j},x_{j}^{+},\widetilde{X}_{j}),$ where $\widetilde{X}_{j}$ is the set of $X_{j}$ perturbed with $x_{i}^{\prime}$. This can be upper-bounded as:
$|\ell_{cont}(x_{j},x_{j}^{+},X_{j})-\ell_{cont}(x_{j},x_{j}^{+},\widetilde{X}_{j})|=\left|\log\frac{\kappa+\operatorname{sim}(x_{j},x_{i})}{\kappa+\operatorname{sim}(x_{j},x_{i}^{\prime})}\right|\leq\log\frac{\kappa+a}{\kappa+b},$
where we define $\kappa=\operatorname{sim}(x_{j},x_{j}^{+})+\sum_{x^{\prime}\in N_{i}}\operatorname{sim}(x_{j},x^{\prime})$ with $N_{i}=X_{i}\!\setminus\!\{x_{i}\}$, $a=e^{1/\tau}$, and $b=e^{-1/\tau}$. Since $a>b$, the function $x\mapsto\frac{x+a}{x+b}$ is strictly decreasing. This can be seen by computing its derivative: $f^{\prime}(x)=\frac{b-a}{(x+b)^{2}}<0$. Therefore, since $\kappa\geq(m-1)e^{-1/\tau}$ and the $\log$ function is increasing, we obtain
$\log\frac{\kappa+a}{\kappa+b}\leq\log\frac{(m-1)e^{-1/\tau}+e^{1/\tau}}{(m-1)e^{-1/\tau}+e^{-1/\tau}}=\log\frac{(m-1)+e^{2/\tau}}{m}.$
Combining these results, we obtain $|\delta^{i}_{\bar{S}_{m}}(\bar{x}_{i})|\leq\frac{4}{\tau}$ and $|\delta^{j}_{\bar{S}_{m}}(\bar{x}_{i})|\leq\log\frac{(m-1)+e^{\frac{2}{\tau}}}{m}$. This gives:
$|\Delta\phi_{\bar{S}}(\bar{x}_{i})|\leq\frac{1}{n}\left(\frac{4}{\tau}+(m-1)\log\frac{(m-1)+e^{\frac{2}{\tau}}}{m}\right)=\frac{C}{n}.$
@@
The rest of the proof of theorem˜3 follows the different steps of the proof presented in McAllester [32], while incorporating the bounded difference assumption. The first step involves showing that we can bound the quantity $\underset{S\sim\mathcal{S}^{n}}{\mathbb{E}}\left[e^{(n-1)h\left(L_{S}(f)-L(f)\right)}\right]$ for a certain non-negative and convex function $h$ (this corresponds to lemmas 5 and 6 in [32]). lemma˜3 below modifies lemma 6 in [32] since we cannot use Hoeffding’s inequality, whereas lemma˜2 is a minor modification of lemma 5 in [32] that allows us to directly apply McDiarmid’s inequality to prove lemma˜3.
@@
The bounded difference assumption allows us to derive the following McDiarmid’s inequality: for $\varepsilon>0$, $\mathbb{P}_{S}\left(|L(f)-L_{S}(f)|\geq\varepsilon\right)\leq 2\exp\left(-\frac{2\varepsilon^{2}}{\sum_{i=1}^{n}c_{i}^{2}}\right)$. Using the change of variable $X=\frac{\sqrt{2}}{C}(L_{S}(f)-L(f))$, we observe that the McDiarmid’s inequality gives: for $x>0$, $\mathbb{P}\left(\frac{\sqrt{2}}{C}|X|\geq x\right)\leq 2\exp\left(-nx^{2}\right)$. Invoking lemma˜2 , we obtain $\mathbb{E}_{S}\left[e^{(n-1)h(X)}\right]\leq 2n$ where $h:x\mapsto\frac{2x^{2}}{C^{2}}$. ∎
Given the bound in lemma˜3, one can derive the PAC-Bayes bound as shown by McAllester [32] (lemma 7–8). We adapt those results in the following statement.
@@
Assume for a fixed $f\sim P$, $\underset{S\sim\mathcal{S}^{n}}{\mathbb{E}}\left[e^{(n-1)h\left(L_{S}(f)-L(f)\right)}\right]\leq 2n$. This implies the bound $\mathbb{E}_{S}\left[\mathbb{E}_{f\sim P}\left[e^{(n-1)h\left(L_{S}(f)-L(f)\right)}\right]\right]\leq 2n$. Applying Markov’s inequality, we obtain that with probability at least $1-\delta$ over $S$, $\mathbb{E}_{f\sim P}\left[e^{(n-1)h\left(L_{S}(f)-L(f)\right)}\right]\leq\frac{2n}{\delta}$. Next, we use a shift of measure: $\mathbb{E}_{f\sim Q}\left[(n-1)h\left(L_{S}(f)-L(f)\right)\right]\leq\mathrm{KL}(Q\|P)+\log\mathbb{E}_{f\sim P}\left[e^{(n-1)h\left(L_{S}(f)-L(f)\right)}\right]$. Combining the previous results, with probability at least $1-\delta$ over dataset $S$:
@@
Rewriting the statement of lemma˜4 shows that, with probability at least $1-\delta$ over i.i.d. dataset $S$, $L(Q)\leq L_{S}(Q)+C\sqrt{\frac{\mathrm{KL}(Q\|P)+\log\frac{2n}{\delta}}{2(n-1)}}$ for all $Q$, completing the proof of theorem˜3.